# Violation of energy conditions and the gravitational radius of the proton

### v2

@@ Abstract
[…]
It is interesting to note that in some holographic models [35] $A=2J$, so (12) holds for $\int\mathrm{d}t\,A(t)\geq 0$ provided the mass $m\geq 0$. In QCD, the integral in (12) is finite because $A-2J\sim 1/t^{3}$ asymptotically [32], but the $-t(A-2J)$ term is not positive definite [12]. Nevertheless, we have confirmed that this inequality indeed holds for the proton EMT when the model GFFs used in our calculations of the point-wise ECs are employed. Thus, assuming that ANEC holds [^2] Besides the explicit proofs in Minkowski spacetime in [29, 30], it is known that violating ANEC would be generally catastrophic, as arguments exist stating that this could be used to build time machines [25, 26] and violate the (generalized) second law of thermodynamics [43]. Finally, we note that in Minkowski spacetime, all null lines are achronal so the ANEC condition discussed here is equivalent to the more stringent achronal averaged null energy condition [27], which is expected to hold even in curved spacetime., (12) provides a new model-independent constraint that should be satisfied by proton GFFs in QCD.
@@
Summary and Discussion: The form factors of the graviton-proton vertex determine the Wigner transform of the QCD energy-momentum tensor in the proton. Its spatial distribution has been used to investigate the mechanical properties of the proton, such as its pressure and shear forces. In this work, motivated by the fact that the EMT is a source term in Einstein’s equations, we have checked its properties according to the classification of Hawking-Ellis using form factors determined from lattice QCD up to momentum transfer ${-t}=2$ GeV2 where we transitioned to the asymptotic forms expected from perturbative QCD. We have found that the EMT in the proton is of “ordinary” type I in the radial tails beyond a few Compton wavelengths where the standard energy conditions such as NEC, WEC, SEC, and DEC are satisfied. In the interior region, however, the proton EMT is of type IV, lacking any causal eigenvectors, and violating all the standard point-wise energy conditions. This unconventional behavior originates from the contribution of the $J$ form factor which determines the $T^{0i}$ component of the EMT. In this regard, the EMT of the pion, where $T^{0i}=0$, will not be of type IV, though the pointwise ECs may still be violated in certain regions.
@@
The exotic properties displayed by the proton EMT, such as the fact that there are regions within the proton where its energy density can vanish or become negative, call for some caution with definitions of the proton’s mechanical properties, such as its mass radius [^3] We refer to ref. [44] for a recent comprehensive review of various proton radius definitions., using this EMT. We propose using the point where the EMT transitions from type IV to type I (thus, necessarily becoming type II at that point) to define a new “gravitational radius” for the proton, which is free from the conceptual issues mentioned above. Using current input from the lattice, we obtain that this occurs at about $1-2$ Compton wavelength. The precise value of this “classicalization” length scale above which typical mechanical properties within the proton are well-defined, standard ECs are fulfilled, and gravity is necessarily attractive, requires determining the gravitational form factors for momentum transfers beyond ${-t}=2$ GeV2, either from lattice QCD or experiments.
@@
Finally, we showed that ANEC translates into the new model-independent constraint (12) for the proton $A(t)$ and $J(t)$ form factors in QCD. A definitive confirmation of ANEC from data reaching to sufficiently high $-t$ could motivate additional theoretical work towards its rigorous proof in asymptotically free theories. In this context, it would be interesting to also check other conditions, such as the quantum null energy condition, which has been thoroughly discussed in the literature [45, 46, 47, 48].
@@ Appendix A Wigner transform of the QCD energy-momentum operator
$x^{0},\vec{x}\left|\,\hat{T}_{\mu\nu}(0)\,\right|x^{0},\vec{x}\right>~{}.$ (18)
@@
$\begin{split}\int_{p}\int_{p^{\prime}}e^{-ip\cdot x}e^{ip^{\prime}\cdot x}\left<p^{\prime}\left|\,\hat{T}_{\mu\nu}(0)\,\right|p\right>&=\int\mathrm{d}^{3}P\int\frac{\mathrm{d}^{3}\Delta}{\sqrt{2p^{0}}\,(2\pi)^{3}\sqrt{2p^{\prime 0}}\,(2\pi)^{3}}\,e^{i\Delta^{0}x^{0}}e^{i(\vec{P}-\frac{1}{2}\vec{\Delta})\cdot\vec{x}}\,e^{-i(\vec{P}+\frac{1}{2}\vec{\Delta})\cdot\vec{x}}\,\left<p^{\prime}\left|\,\hat{T}_{\mu\nu}(0)\,\right|p\right>~{}.\end{split}$ (19)
@@
and interpret $x^{0},\vec{x}\left.\right|x^{0},\vec{x}\right>$ in the sense of distributions to see that its inverse may be represented as $(2\pi)^{3}\delta(\vec{P}-\vec{Q})$ when working with momentum space basis vectors. Thus, we find that
$x^{0},\vec{x}\left|\,\hat{T}_{\mu\nu}(0)\,\right|x^{0},\vec{x}\right>=\left<x^{0},\vec{x}\left.\right|x^{0},\vec{x}\right>\int\frac{\mathrm{d}^{3}\Delta}{(2\pi)^{3}}\,\frac{1}{\sqrt{4P_{0}^{2}-\Delta_{0}^{2}}}\,e^{i\Delta\cdot x}\,\left<\vec{P}+\frac{\vec{\Delta}}{2},s\left|\,\hat{T}_{\mu\nu}(0)\,\right|\vec{P}-\frac{\vec{\Delta}}{2},s\right>~{},$ (23)
@@
The static EMT (3) of Polyakov et al. is recovered by taking $\vec{P}=0$, the “Breit frame”. Eq. (24) represents the Wigner-Weyl transform of $\hat{T}_{\mu\nu}(0)$ to the phase-space quasi-probability distribution $T_{\mu\nu}(x^{0},\vec{x},\vec{P})$. Note that this is conserved,
@@ Appendix B Components of the Wigner EMT in terms of the form factors
$\frac{\vec{\Delta}}{2},s\left|\,\hat{T}_{00}(0)\,\right|-\frac{\vec{\Delta}}{2},s\right>=2mE\left[A(t)+\frac{-t}{4m^{2}}\left(A(t)-2J(t)+D(t)\right)\right]~{}.$ (28)
@@
[…]
@@
$p^{\prime},s\left|\,\hat{T}^{0i}(0)\,\right|p,s\right>=-E\,J(t)\,i\,(\vec{\Delta}\times\vec{\sigma}_{ss})^{i}~{}.$ (33)
@@
Here, $\sigma_{ss}^{i}=\xi^{\dagger}_{s}\sigma^{i}\xi_{s}$. For proton spin $+\frac{1}{2}$ along the $z$-axis (i.e. $\theta=0$ so that $\hat{n}=\hat{z}$, and $\xi_{s}=\xi_{\uparrow}$) we have $\vec{\sigma}_{ss}=(0,0,1)$ and $(\vec{\Delta}\times\vec{\sigma}_{ss})^{x}=\Delta^{y}$, $(\vec{\Delta}\times\vec{\sigma}_{ss})^{y}=-\Delta^{x}$, $(\vec{\Delta}\times\vec{\sigma}_{ss})^{z}=0$. Then, from Eq. (3),
@@
Specifically for spin $+\frac{1}{2}$ along the $z$-axis,
@@ Appendix C Derivation of the eigenvalues of $T_{\mu\nu}$
In this section, we adhere to the common mostly plus choice for the metric tensor $g_{\mu\nu}$ used in the GR literature [17]. The Minkowski metric in Cartesian coordinates is now $\eta_{\mu\nu}=\mathrm{diag}(-1,1,1,1)$.
@@
As mentioned in the main text, the energy conditions we consider are the null energy condition (NEC), the weak energy condition (WEC), the dominant energy condition (DEC), and the strong energy condition (SEC), whose definitions were given in the main text, and are also be found in [17].
@@ Appendix D Details about the ANEC calculation
$\begin{split}\bar{u}(p^{\prime})u(p)&={\cal N}^{-1}\left[2\left(P_{0}^{2}-\vec{P}^{2}+P_{0}m\right)+i\epsilon^{0ijk}P_{i}\Delta_{j}S_{k}\right]={\cal N}^{-1}\left[2\left(P_{0}^{2}-\vec{P}^{2}+P_{0}m\right)-i\vec{P}\cdot(\vec{\Delta}\times\vec{S})\right]~{},\\ \bar{u}(p^{\prime})i\sigma^{\mu\rho}\Delta_{\rho}u(p)&={\cal N}^{-1}\left\{P^{\mu}\Delta^{2}+m(\eta^{\mu 0}\Delta^{2}-\Delta^{\mu}\Delta^{0})+2\left[(P^{0}+m)i\epsilon^{\mu\nu\rho\lambda}\Delta_{\nu}P_{\rho}S_{\lambda}-\frac{\Delta^{2}}{4}i\epsilon^{\mu\nu\rho 0}\Delta_{\nu}S_{\rho}-(P\cdot S)i\epsilon^{\mu\nu\rho 0}\Delta_{\nu}P_{\rho}\right]\right\}~{}\end{split}$ (59)

### v3

@@ V New constraints on GFFs from ANEC
VI Summary and Discussion
@@
The form factors of the graviton-proton vertex determine the Wigner transform of the QCD energy-momentum tensor in the proton. Its spatial distribution has been used to investigate the mechanical properties of the proton, such as its pressure and shear forces. In this work, motivated by the fact that the EMT is a source term in Einstein’s equations, we have checked its properties according to the classification of Hawking-Ellis using form factors determined from lattice QCD up to momentum transfer t=2{-t}=2 GeV2 where we transitioned to the asymptotic forms expected from perturbative QCD. We have found that the EMT in the proton is of “ordinary” type I in the radial tails beyond a few Compton wavelengths where the standard energy conditions such as NEC, WEC, SEC, and DEC are satisfied. In the interior region, however, the proton EMT may be of type IV, lacking any causal eigenvectors, and violating all the standard point-wise energy conditions. This unconventional behavior originates from the contribution of the $J$ form factor which determines the $T^{0i}$ component of the EMT. In this regard, the EMT of the pion, where $T^{0i}=0$, will not be of type IV, though the pointwise ECs may still be violated in certain regions.
@@
The exotic properties potentially displayed by the proton EMT, such as the fact that there may be regions within the proton where its energy density can vanish or become negative, call for some caution with definitions of the proton’s mechanical properties, such as its mass radius [^4] We refer to Ref. [49] for a recent comprehensive review of various proton radius definitions., using this EMT. We propose using the point where the EMT transitions from type IV to type I (thus, necessarily becoming type II at that point) to define a new “gravitational radius” for the proton, which is free from the conceptual issues mentioned above. Using current input from the lattice, we obtain that this occurs at about 121-2 Compton wavelength. The precise value of this “classicalization” length scale above which typical mechanical properties within the proton are well-defined, standard ECs are fulfilled, and gravity is necessarily attractive, requires determining the gravitational form factors for momentum transfers beyond t=2{-t}=2 GeV2, either from lattice QCD or experiments.
@@
Finally, we showed that ANEC translates into the new model-independent, non-perturbative constraints (12) and (13) for the proton $A(t)$ and $J(t)$ form factors and the $A(t)$ pion form factor in QCD, respectively. A definitive confirmation of ANEC from data reaching to sufficiently high t-t could motivate additional theoretical work towards new rigorous proofs in asymptotically free theories [^5] ANEC was proven to hold in unitary, Lorentz invariant quantum field theories in Minkowski spacetime in [30] using quantum information-theoretic inequalities, such as the monotonicity of the relative entropy. In [31], a proof was obtained using causality considerations.. In this context, it would be interesting to also check other conditions, such as the quantum null energy condition, which has been thoroughly discussed in the literature [50, 51, 52, 53].
@@ Appendix A Wigner transform of the QCD energy-momentum operator
$\left<x^{0},\vec{x}\left|\,\hat{T}_{\mu\nu}(0)\,\right|x^{0},\vec{x}\right$>~. (19)
@@
[…]
@@
and interpret $\left<x^{0},\vec{x}\left.\right|x^{0},\vec{x}\right$> in the sense of distributions to see that its inverse may be represented as $(2\pi)^{3}\delta(\vec{P}-\vec{Q})$ when working with momentum space basis vectors. Thus, we find that
$\left<x^{0},\vec{x}\left|\,\hat{T}_{\mu\nu}(0)\,\right|x^{0},\vec{x}\right>=\left<x^{0},\vec{x}\left.\right|x^{0},\vec{x}\right>\int\frac{\mathrm{d}^{3}\Delta}{(2\pi)^{3}}\,\frac{1}{\sqrt{4P_{0}^{2}-\Delta_{0}^{2}}}\,e^{i\Delta\cdot x}\,\left<\vec{P}+\frac{\vec{\Delta}}{2},s\left|\,\hat{T}_{\mu\nu}(0)\,\right|\vec{P}-\frac{\vec{\Delta}}{2},s\right$>~, (24)
@@
The static EMT of Polyakov et al. is recovered by taking $\vec{P}=0$, the “Breit frame”, which gives the energy-momentum tensor in Eq. (3) of the main text:
$T_{\mu\nu}(\vec{x}\,)=\int\frac{\mathrm{d}^{3}\Delta}{2E\,(2\pi)^{3}}\,e^{-i\vec{\Delta}\cdot\vec{x}}\,\left<\frac{\vec{\Delta}}{2},s\left|\,\hat{T}_{\mu\nu}(0)\,\right|-\frac{\vec{\Delta}}{2},s\right$>. (26)
Eq. (25) represents the Wigner-Weyl transform of $\hat{T}_{\mu\nu}(0)$ to the phase-space quasi-probability distribution $T_{\mu\nu}(x^{0},\vec{x},\vec{P})$. Note that this is conserved,
@@ Appendix B Components of the Wigner EMT in terms of the form factors
$\left<\frac{\vec{\Delta}}{2},s\left|\,\hat{T}_{00}(0)\,\right|-\frac{\vec{\Delta}}{2},s\right>=2mE\left[A(t)+\frac{-t}{4m^{2}}\left(A(t)-2J(t)+D(t)\right)\right]~.$ (30)
@@
The second term on the r.h.s. immediately gives $2E^{2}/m$. For the first term on the r.h.s., using the chiral/Weyl representation of $\gamma^{0}$ [34], we have $\gamma^{0}u_{s}(p)=u_{s}(p^{\prime})$ since $\vec{p}=-\vec{p}^{\,\prime}$; so, for that term one obtains $\bar{u}_{s}(p^{\prime})u_{s}(p^{\prime})=2m$. In all, the r.h.s. evaluates to $2\vec{p}^{\,2}/m=-(1/2)\vec{\Delta}^{2}/m=t/2m$. Hence,
@@
$\left<p^{\prime},s\left|\,\hat{T}^{0i}(0)\,\right|p,s\right>=-E\,J(t)\,i\,(\vec{\Delta}\times\vec{\sigma}_{ss})^{i}~.$ (35)
@@
Here, $\sigma_{ss}^{i}=\xi^{\dagger}_{s}\sigma^{i}\xi_{s}$. For proton spin $\frac{1}{2}$ along the $z$-axis (i.e. $\theta=0$ so that $\hat{n}=\hat{z}$, and $\xi_{s}=\xi_{\uparrow}$) we have $\vec{\sigma}_{ss}=(0,0,1)$ and $(\vec{\Delta}\times\vec{\sigma}_{ss})^{x}=\Delta^{y}$, $(\vec{\Delta}\times\vec{\sigma}_{ss})^{y}=-\Delta^{x}$, $(\vec{\Delta}\times\vec{\sigma}_{ss})^{z}=0$. Then, from Eq. (3),
@@
Specifically for spin $\frac{1}{2}$ along the $z$-axis,
@@ Appendix C Derivation of the eigenvalues of $T_{\mu\nu}$
In this appendix, we adhere to the common mostly plus choice for the metric tensor $g_{\mu\nu}$ used in the GR literature [18]. The Minkowski metric in Cartesian coordinates is now $\eta_{\mu\nu}=\mathrm{diag}(-1,1,1,1)$.
@@
As mentioned in the main text, the energy conditions we consider are the null energy condition, the weak energy condition, the dominant energy condition, and the strong energy condition, whose definitions were given in the main text, and are also be found in [18].
@@ Appendix D Details about the ANEC calculation
We first consider the averaged null energy condition for the spin-$1/2$ proton and then for the spin-0 pion.
@@
[…]
@@
We now turn to a spin-0 pion. The matrix elements of its EMT are
$\begin{split}\left<\vec{P}+\frac{\vec{\Delta}}{2}\left|\,\hat{T}^{00}(0)\,\right|\vec{P}-\frac{\vec{\Delta}}{2}\right>&=2(P^{0})^{2}A(t)+\frac{1}{2}\vec{\Delta}^{2}D(t)~,\\ \left<\vec{P}+\frac{\vec{\Delta}}{2}\left|\,\hat{T}^{0i}(0)\,\right|\vec{P}-\frac{\vec{\Delta}}{2}\right>&=2P^{0}P^{i}A(t)+\frac{1}{2}\Delta^{0}\Delta^{i}D(t)~,\\ \left<\vec{P}+\frac{\vec{\Delta}}{2}\left|\,\hat{T}^{ij}(0)\,\right|\vec{P}-\frac{\vec{\Delta}}{2}\right>&=2P^{i}P^{j}A(t)+\frac{1}{2}\left(\Delta^{i}\Delta^{j}+\delta^{ij}\Delta^{2}\right)D(t)~,\end{split}$ (66)
These are the expressions to be used in eq. (25) for $T_{\mu\nu}(x^{0},\vec{x},\vec{P})$.
The GFFs of a spin 0 particle have monopole form,
$A(t)=\frac{\alpha_{Aq}}{1-t/\Lambda_{Aq}^{2}}+\frac{\alpha_{Ag}}{1-t/\Lambda_{Ag}^{2}}~,$ (67)
and similar for $D(t)$. The monopole masses $\Lambda_{Aq},\Lambda_{Ag}$ in the quark and gluon contributions have been determined on the lattice in ref. [55], again up to t=2{-t}=2 GeV2. In the asymptotic regime the pion GFFs from perturbative QCD have been found to scale $\sim 1/t$ [47, 48].
From eq. (25) we obtain the components of the Wigner transform; for simplicity, we again choose the Breit frame. We then have
$\begin{split}T^{00}(\vec{x})&=\frac{[\Delta^{2}A(-\Delta^{2})]_{\infty}+[\Delta^{2}D(-\Delta^{2})]_{\infty}}{4\pi^{2}\,r^{2}}\\ &+\int\frac{\mathrm{d}^{3}\Delta}{(2\pi)^{3}}\,\,e^{-i\vec{\Delta}\cdot\vec{x}}\,\left[EA(t)+\frac{1}{4E}\vec{\Delta}^{2}D(t)-\frac{[\Delta^{2}A(-\Delta^{2})]_{\infty}}{2\Delta}-\frac{[\Delta^{2}D(-\Delta^{2})]_{\infty}}{2\Delta}\right]~.\end{split}$ (68)
Here, $\Delta^{2}A(-\Delta^{2})]_{\infty}=\lim_{\Delta\to\infty}\,\Delta^{2}A(-\Delta^{2})$ with $\Delta=|\vec{\Delta}|$. In the Breit frame, $T^{0i}=0$. Lastly,
$\begin{split}T^{ij}(\vec{x})&=-[\Delta^{2}D(-\Delta^{2})]_{\infty}\frac{x^{i}x^{j}}{2\pi^{2}r^{4}}+\int\frac{\mathrm{d}^{3}\Delta}{(2\pi)^{3}}\,\,e^{-i\vec{\Delta}\cdot\vec{x}}\,\left(\Delta^{i}\Delta^{j}-\delta^{ij}\vec{\Delta}^{2}\right)\left[\frac{D(t)}{4E}-\frac{[\Delta^{2}D(-\Delta^{2})]_{\infty}}{2\Delta^{3}}\right]~.\end{split}$ (69)
[…]
$\begin{split}\int\limits_{-\infty}^{\infty}\mathrm{d}\zeta\,\ell^{\mu}\ell^{\nu}\,T_{\mu\nu}(\vec{x})&=\int\limits_{-\infty}^{\infty}\mathrm{d}\zeta\,\left[T_{00}(\vec{x})+\hat{n}^{i}\hat{n}^{j}T_{ij}(\vec{x})\right]~.\end{split}$ (70)
The first term with $T_{00}$ gives
$\begin{split}&\frac{[\Delta^{2}A(-\Delta^{2})]_{\infty}+[\Delta^{2}D(-\Delta^{2})]_{\infty}}{4\pi L\sqrt{{x_{0}}^{2}-(\vec{x}_{0}\cdot\hat{n})^{2}}}\\ &+\int\frac{\mathrm{d}^{2}\Delta_{\perp}}{(2\pi)^{2}L}\,e^{-i\vec{\Delta}_{\perp}\cdot\vec{x}_{0\perp}}\left[EA(-\Delta_{\perp}^{2})+\frac{\Delta_{\perp}^{2}}{4E}D(-\Delta_{\perp}^{2})-\frac{[\Delta^{2}A(-\Delta^{2})]_{\infty}}{2\Delta_{\perp}}-\frac{[\Delta^{2}D(-\Delta^{2})]_{\infty}}{2\Delta_{\perp}}\right]~,\end{split}$ (71)
while the term involving $T_{ij}$ gives
$\begin{split}&-\frac{[\Delta^{2}D(-\Delta^{2})]_{\infty}}{4\pi L\sqrt{{x_{0}}^{2}-(\vec{x}_{0}\cdot\hat{n})^{2}}}-\int\frac{\mathrm{d}^{2}\Delta_{\perp}}{(2\pi)^{2}L}\,\,e^{-i\vec{\Delta}_{\perp}\cdot\vec{x}_{0\perp}}\,\vec{\Delta}_{\perp}^{2}\left[\frac{D(-\Delta_{\perp}^{2})}{4E}-\frac{[\Delta^{2}D(-\Delta^{2})]_{\infty}}{2\Delta_{\perp}^{3}}\right]\end{split}$ (72)
Adding both contributions,
$\begin{split}\mathrm{ANEC}_{\pi}&=\frac{[\Delta^{2}A(-\Delta^{2})]_{\infty}}{4\pi L\,|\vec{x}_{0\perp}|}+\int\frac{\mathrm{d}^{2}\Delta_{\perp}}{(2\pi)^{2}L}\,e^{-i\vec{\Delta}_{\perp}\cdot\vec{x}_{0\perp}}\left[EA(-\Delta_{\perp}^{2})-\frac{[\Delta^{2}A(-\Delta^{2})]_{\infty}}{2\Delta_{\perp}}\right]~.\end{split}$ (73)
[…]